OpenStudy (anonymous):

What is the difference of 2 over the quantity x minus 4 − the quantity 7 times x minus 28, end quantity over the quantity x minus 4, end quantity, squared?

OpenStudy (anonymous):

\[\left(\begin{matrix}2 \\x-4 \end{matrix}\right) - \left(\begin{matrix}7x-28 \\(x-4) ^{2}\end{matrix}\right)\]

OpenStudy (pfenn1):

You mean?\[\frac{2}{x-4}- \frac{7x-28 }{(x-4) ^{2}}\]

OpenStudy (oaktree):

-5 / (x-4), or \[-5\div(x-4)\]

OpenStudy (pfenn1):

@OakTree, help to see how to get the answer, don't just provide the answer. Thanks.

OpenStudy (anonymous):

thanks

OpenStudy (pfenn1):

You're welcome.